What mass of gold (in grams) is plated by a 25 minute flow of 5.5 A current?

Keeping in mind that gold can be plated out of solution containing Au3+ according to the half reaction.

Au3+(aq) + 3 e- → Au (s)

(hint: your answer should not contain units and have two significant figures)
1
Expert's answer
2017-05-05T07:01:38-0400
m(Au)=(M(Au)It)/nF=(197g/mol∙5.5A∙25min∙60sec/min)/(3∙96500C/mol)=5.6